OpenStudy (saifoo.khan):

The expression x3 + ax2 – 15x + b has a factor x – 2 and leaves a remainder of 75 when divided by x + 3. Find the value of a and of b.

OpenStudy (saifoo.khan):

c'mon guyz! :s

OpenStudy (anonymous):

a is -8 and b is 54

OpenStudy (saifoo.khan):

a = 7 b = -6

OpenStudy (anonymous):

This polynomial has a factor of (x-2), so plug in x=2 to x3 + ax2 – 15x + b =0, you'll get 1 equation

OpenStudy (anonymous):

When you divide x3 + ax2 – 15x + b by (x+3), you'll get 75; so based on this info, plug in x=-3 to x3 + ax2 – 15x + b =75, you get the 2nd equation. Now you have 2 variables with 2 equations. Nice problem.

OpenStudy (anonymous):

The 2 equations are: 4a + b = 22 and 9a + b = 57. As you pointed out, the solutions are a=7 and b=-6

OpenStudy (anonymous):

Hey saifoo, yr ans is correct I rechecked and got same values for a and b